# Commutation relations for two-sided radial SLE

### v2

@@ 2.1 Definition of locally commuting 2-radial SLE
Let $D\subset\mathbb{D}$ be a simply connected domain containing 00 and let $x_{1},x_{2}$ be distinct prime ends of $D$. Let $U_{1},U_{2}$ be, respectively, closed neighborhoods of $x_{1}$ and $x_{2}$ in $D$ that do not contain 00 and such that $U_{1}\cap U_{2}=\emptyset$. We will consider probability measures $\mathbb{P}_{(D;\,x_{1},x_{2})}^{(U_{1},U_{2})}$ on pairs of unparametrized continuous curves in $U_{1}$ and $U_{2}$ starting from $x_{1}$ and $x_{2}$, and exiting $U_{1}$ and $U_{2}$ almost surely. We call that such a family of measures indexed by different choices of $(U_{1},U_{2})$ compatible if for all $U_{1}\subset U_{1}^{\prime}$ and $U_{2}\subset U_{2}^{\prime}$, we have $\mathbb{P}_{(D;\,x_{1},x_{2})}^{(U_{1},U_{2})}$ is obtained from restricting the curves under $\mathbb{P}_{(D;\,x_{1},x_{2})}^{(U_{1}^{\prime},U_{2}^{\prime})}$ to the part before first exiting the subdomains $U_{1}$ and $U_{2}$.
@@
    Conformal invariance: If $\varphi:D\to D^{\prime}$ is a conformal map fixing 00, then the pullback measure
@@ 2.2 Commutation relations
where $\delta=\varepsilon\left(\sin((\theta_{2}-\theta_{1})/2)\right)^{-2}$ so that $\varepsilon^{\prime}=\varepsilon(1-\delta)$.
@@
We note that $\mathcal{Z}$ does not always descend to a function on $(\mathbb{R}/2\pi\mathbb{Z})^{2}$.
@@
$\partial_{1}b_{2}=\partial_{2}b_{1},$ (2.9)
@@
$2\displaystyle 0=\partial_{1}\left(\frac{\partial_{j}\mathcal{Z}}{\mathcal{Z}}\right)+\partial_{2}\left(\frac{\partial_{j}\mathcal{Z}}{\mathcal{Z}}\right),\qquad\forall j=1,$ 2 . (2.16)
@@ 2.3 Solutions to commutation relations with interchangeability
Let us give a comment on the $C^{2}$ requirement on $b_{1},b_{2}$ in the assumption (MARG). We assume such $C^{2}$ requirement as part of the axioms at the beginning. But in fact, this is not necessary. Suppose we relax the requirement and only assume $b_{1},b_{2}$ in (MARG) are continuous, we are still able to derive (2.14) and (2.15) as weak solutions. The operators in the left-hand side of radial BPZ equations (2.14) and (2.15) are hypoelliptic, due to a general characterization by Hörmander [15], see also [10, Lemma 5] and [30, Sect. 2.3.3]. Therefore, the weak solutions are strong solutions which are in fact $C^{\infty}$, and consequently $b_{1},b_{2}$ are $C^{\infty}$. Note that such analysis does not work for $\kappa=0$ since the corresponding BPZ equation (4.1) is nonlinear.
@@ 3.1 Radial SLE
Lemma 3.1 (See [16, Prop. 5] or [14, Prop. 2.2]).
@@
It is explained in the proof of [16, Prop. 5] that the term $m_{t}=m_{\mathbb{D}}(\eta_{[0,t]},K)$ is the same as the Brownian loop measure of loops that intersect both $\eta_{[0,t]}$ and $K$ when $\eta_{[0,t]}\cap K=\emptyset$.
@@
This is proved in [28, Prop. 3.30] and [28, Sect. 4]. See also [20]. ∎
@@ 3.2 Two-sided radial SLE with spiral
In this section, we introduce two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral when $\kappa\in(0,8)$ by reweighting two independent radial SLEκ. The two-sided radial SLE analyzed in [20, 13, 23, 14] is a special case where the spiraling rate $\mu=0$.
@@
×$\times\mathcal{G}_{\mu}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)\exp\left(\frac{c}{2}m_{\mathbf{t}}\right),$ (3.3)
@@
Let us first compute the variations of terms appearing in (3.5). From standard calculations (see e.g. [14, Lem. 3.2]), we have the following variational formula of the capacity parameterizations:
@@
When $\kappa\leq 4$, two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral is well-defined for all time and the two curves do not touch each other before they reach the origin. When $\kappa\in(4,8)$, the above definition for two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral is only defined up to the times the two curves touch each other. When the spiraling rate is 00, we obtain the standard two-sided radial $\operatorname{SLE}$ analyzed in [20, 13, 23, 14].
@@
Remark 3.8.
Two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral can be generalized to the multiple-sided case: $N$-sided radial $\operatorname{SLE}_{\kappa}$ with spiraling rate $\mu$ for $N\geq 2$ can be defined a similar way as in Definition 3.6 where the partition function $\mathcal{G}_{\mu}$ shall be replaced by
$\mathcal{G}_{\mu}(\theta_{1},\ldots,\theta_{N})=\prod_{1\leq i<j\leq N}\left(\sin((\theta_{j}-\theta_{i})/2)\right)^{2/\kappa}\times\exp\left(\frac{\mu}{\kappa}\sum_{j=1}^{N}\theta_{j}\right),$
for $\theta_{1}<\cdots<\theta_{N}<\theta_{1}+2\pi$. When $\mu=0$, it is the same as the partition function for $N$-sided radial $\operatorname{SLE}_{\kappa}$ in [14].
@@ 3.3 Resampling property of two-sided radial SLE with spiral
  • Given $\eta^{(1)}$, the conditional law of $\eta^{(2)}$ is chordal $\operatorname{SLE}_{\kappa}$ in $\mathbb{D}\setminus\eta^{(1)}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to 00.
@@
Radial SLE with spiral appears as a flow line in the setup of imaginary geometry [28]. Furthermore, two-sided radial $\operatorname{SLE}_{\kappa}$ with spiraling rate $\mu$ can be viewed as a pair of flow lines of
@@
where $\Gamma$ is a GFF in $\mathbb{D}$ with properly chosen boundary data, and the angles of the two flow lines are also chosen properly. See Figure 2. Using such coupling, one is able to derive the resampling property in Theorem 3.9, see [28, Prop. 3.28]. However, our proof of the resampling property in Section 3.2 does not use the coupling with imaginary geometry. We derive it directly using a refined analysis of the Radon–Nikodym derivative $M_{\mathbf{t}}(\mathcal{G}_{\mu})$ in Definition 3.6.
@@
$\frac{M_{(t_{1},t_{2})}(\mathcal{G}_{\mu})}{M_{(t_{1},0)}(\mathcal{G}_{\mu})}=$ $\underbrace{\frac{\phi_{\mathbf{t},2}^{\prime}\left(\xi_{t_{2}}^{(2)}\right)^{h}g_{\mathbf{t},2}^{\prime}(0)^{\tilde{h}}\exp\left(\frac{c}{2}m_{\mathbf{t}}\right)}{\left(\phi_{t_{1}}^{(1)}\right)^{\prime}(\theta_{2})^{h}\left(g_{t_{1}}^{(1)}\right)^{\prime}(0)^{\tilde{h}}}}_{=:P_{\mathbf{t}}}\times\underbrace{g_{\mathbf{t},1}^{\prime}(0)^{\frac{3-\mu^{2}}{2\kappa}}\frac{\mathcal{G}_{\mu}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)}{\mathcal{G}_{\mu}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)}\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{h}}_{=:R_{\mathbf{t}}}.$
@@
$R_{\mathbf{t}}=g_{\mathbf{t},1}^{\prime}(0)^{\frac{3-\mu^{2}}{2\kappa}}\frac{\mathcal{G}_{\mu}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)}{\mathcal{G}_{\mu}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)}\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{h}.$ (3.16)
Combining Lemma 3.3 and Lemma 3.11, we see that radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ converges to chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)};\mathrm{e}^{\mathfrak{i}\theta_{2}},0)$ as $t_{1}\to\infty$; and we will show in Lemma 3.12 that $R_{\mathbf{t}}\to 1$ almost surely as $t_{1}\to\infty$. Combining these two parts, the conditional law of $\eta^{(2)}_{[0,t_{2}]}$ given $\eta^{(1)}$ is the same as chordal $\operatorname{SLE}_{\kappa}$ as desired. ∎
@@
Fix $\kappa\in(0,4]$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. We assume $\eta^{(1)}\in\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}};0)$, namely a simple curve in $\mathbb{D}$ from $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ to 00.
  • For $t_{1}\in(0,\infty)$, we denote by $Q_{t_{1}}$ the law of a radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$.
  • We denote by $Q_{\infty}$ the law of a chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)};\mathrm{e}^{\mathfrak{i}\theta_{2}},0)$.
Then for any $t_{1},t_{2}\in(0,\infty)$, the law of the curve $\eta^{(2)}$ restricted to $[0,t_{2}]$ (under intrinsic capacity parametrization) under $Q_{t_{1}}$ and $Q_{\infty}$ are absolutely continuous. We have
$\lim_{t_{1}\to\infty}\frac{\mathrm{d}Q_{t_{1}}}{\mathrm{d}Q_{\infty}}\left(\eta^{(2)}_{[0,t_{2}]}\right)=1,\quad Q_{\infty}-a.s.$
@@
It is proved in [17, Lem. 3.2] and [21]. To be self-contained, we include its short proof adapted to our setting in Appendix B. ∎
@@
We first argue that, almost surely, the difference
$\Delta_{t_{1}}:=\phi_{t_{1}}^{(1)}(\theta_{2})-\xi_{t_{1}}^{(1)}$
is bounded away from 00 and $2\pi$. As the marginal law of $\eta^{(1)}$ is radial $\operatorname{SLE}_{\kappa}^{\mu}(2)$, we have
$\mathrm{d}\Delta_{t_{1}}=-\sqrt{\kappa}\mathrm{d}B_{t_{1}}^{(1)}+2\cot(\Delta_{t_{1}}/2)\mathrm{d}t_{1}-\mu\mathrm{d}t_{1},$
where $B^{(1)}$ is one-dimensional Brownian motion. Roughly speaking, when $\Delta_{t_{1}}$ is close to zero, it is absolutely continuous with respect to the Bessel process of dimension $8/\kappa+1\geq 3$; this explains that it is bounded away from zero. We will give more precise details below. Define
$f(\theta)=\int_{\theta}^{2\pi-\theta}\exp(2\mu u/\kappa)\left(\sin(u/2)\right)^{-8/\kappa}\mathrm{d}u,\quad\text{for }\theta\in(0,2\pi),$
then $f(\Delta_{t})$ is a local martingale. Suppose $\Delta_{0}\in(0,2\pi)$. For $n\geq 1$, define $T_{n}=\inf\{t:\Delta_{t}=2^{-n}\text{ or }\Delta_{t}=2\pi-2^{-n}\}$. For $n$ large enough, we have $\Delta_{0}\in(2^{-n},2\pi-2^{-n})$. Optional stopping theorem gives $\mathbb{E}[f(\Delta_{T_{n}})]=f(\Delta_{0})$. Thus
$\mathbb{P}[T_{n}<\infty]\sim 2^{-n(8/\kappa-1)},\quad\text{for large $n$}.$
In particular, we have $\sum_{n}\mathbb{P}[T_{n}<\infty]<\infty$, and Borel-Cantelli lemma tells that almost surely, there exists $n_{0}$ such that $T_{n_{0}}=\infty$. In other words, almost surely, there exists $n_{0}$ such that $\Delta_{t}\in(2^{-n_{0}},2\pi-2^{-n_{0}})$ for all $t$.
Next, we show that $R_{\mathbf{t}}\to 1$ as $t_{1}\to\infty$. We write $I_{t_{1}}$ for the arc in $\partial\mathbb{D}$ that is the image of both sides of $\eta^{(1)}_{[0,t_{1}]}$ under the conformal map $g_{t_{1}}^{(1)}$ extended to the boundary, see Figure 5. It is easy to see the harmonic measure of $\partial\mathbb{D}$ seen from 00 of the domain $\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}$ is decreasing to 00 as $t_{1}\to\infty$. Therefore,

### v3

@@ 2.1 Definition of locally commuting 2-radial SLE
Let $D\subset\mathbb{D}$ be a simply connected domain containing 0 and let $x_{1},x_{2}$ be distinct prime ends of $D$. Let $U_{1},U_{2}$ be, respectively, closed neighborhoods of $x_{1}$ and $x_{2}$ in $D$ that do not contain 0 and such that $U_{1}\cap U_{2}=\emptyset$. We will consider probability measures $\mathbb{P}_{(D;\,x_{1},x_{2})}^{(U_{1},U_{2})}$ on pairs of unparametrized continuous curves in $U_{1}$ and $U_{2}$ starting from $x_{1}$ and $x_{2}$, and exiting $U_{1}$ and $U_{2}$ almost surely. We call that such a family of measures indexed by different choices of $(U_{1},U_{2})$ compatible if for all $U_{1}\subset U_{1}^{\prime}$ and $U_{2}\subset U_{2}^{\prime}$, we have $\mathbb{P}_{(D;\,x_{1},x_{2})}^{(U_{1},U_{2})}$ is obtained from restricting the curves under $\mathbb{P}_{(D;\,x_{1},x_{2})}^{(U_{1}^{\prime},U_{2}^{\prime})}$ to the part before first exiting the subdomains $U_{1}$ and $U_{2}$.
@@
    Conformal invariance: If $\varphi:D\to D^{\prime}$ is a conformal map fixing 0, then the pullback measure
@@ 2.2 Commutation relations
where $\delta=\varepsilon\left(\sin((\theta_{2}-\theta_{1})/2)\right)^{-2}$ so that $\varepsilon^{\prime}=\varepsilon(1-\delta)+o(\varepsilon^{2})$.
@@
We note that $\mathcal{Z}$ does not always descend to a function on $(\mathbb{R}/2\pi\mathbb{Z})^{2}$. The radial BPZ equations (or BPZ–Cardy equations) were also derived in the context of conformal field theory in [23, 24].
@@
$\kappa\partial_{1}b_{2}=\kappa\partial_{2}b_{1},$ (2.9)
@@
$0=\partial_{1}\left(\frac{\partial_{j}\mathcal{Z}}{\mathcal{Z}}\right)+\partial_{2}\left(\frac{\partial_{j}\mathcal{Z}}{\mathcal{Z}}\right),\qquad\forall j=1,2.$ (2.16)
@@ 2.3 Solutions to commutation relations with interchangeability
Let us give a comment on the $C^{2}$ requirement on $b_{1},b_{2}$ in the assumption (MARG). We assume such $C^{2}$ requirement as part of the axioms at the beginning. But in fact, this is not necessary. Suppose we relax the requirement and only assume $b_{1},b_{2}$ in (MARG) are continuous, we are still able to derive (2.14) and (2.15) as weak solutions. The operators in the left-hand side of radial BPZ equations (2.14) and (2.15) are hypoelliptic, due to a general characterization by Hörmander [18], see also [13, Lemma 5] and [35, Sect. 2.3.3]. Therefore, the weak solutions are strong solutions which are in fact $C^{\infty}$, and consequently $b_{1},b_{2}$ are $C^{\infty}$. Note that such analysis does not work for $\kappa=0$ since the corresponding BPZ equation (4.1) is nonlinear.
@@ 3.1 Radial SLE
Lemma 3.1 (See [20, Prop. 5] or [17, Prop. 2.2]).
@@
It is explained in the proof of [20, Prop. 5] that the term $m_{t}=m_{\mathbb{D}}(\eta_{[0,t]},K)$ is the same as the Brownian loop measure of loops that intersect both $\eta_{[0,t]}$ and $K$ when $\eta_{[0,t]}\cap K=\emptyset$.
@@
This is proved in [33, Prop. 3.30] and [33, Sect. 4]. See also [25]. ∎
@@ 3.2 Two-sided radial SLE with spiral
In this section, we define two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral when $\kappa\in(0,8)$ by reweighting two independent radial SLEκ. The two-sided radial SLE analyzed in [25, 16, 28, 17] is a special case where the spiraling rate $\mu=0$.
@@
$\times\mathcal{G}_{\mu}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)\exp\left(\frac{c}{2}m_{\mathbf{t}}\right),$ (3.3)
@@
Let us first compute the variations of terms appearing in (3.5). From standard calculations (see e.g. [17, Lem. 3.2]), we have the following variational formula of the capacity parameterizations:
@@
From Corollary 3.7 below, it follows that when $\kappa\leq 4$, two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral is well-defined for all time and the two curves do not touch each other before they reach the origin. When $\kappa\in(4,8)$, the above definition for two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral is only defined up to the times the two curves touch each other. When the spiraling rate is 0, we obtain the standard two-sided radial $\operatorname{SLE}$ analyzed in [25, 16, 28, 17].
@@ 3.3 Resampling property of two-sided radial SLE with spiral
  • Given $\eta^{(1)}$, the conditional law of $\eta^{(2)}$ is chordal $\operatorname{SLE}_{\kappa}$ in $\mathbb{D}\setminus\eta^{(1)}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to 0.
@@
Radial SLE with spiral appears as a flow line in the setup of imaginary geometry [33]. An alternative way to construct a two-sided radial $\operatorname{SLE}_{\kappa}$ with spiraling rate $\mu$ is to take pair of flow lines of
@@
where $\Gamma$ is a GFF in $\mathbb{D}$ with properly chosen boundary data, and the angles of the two flow lines are also chosen properly. See Figure 2. Using such coupling, one is able to derive the resampling property in Theorem 3.8, see [33, Prop. 3.28]. However, our proof of the resampling property in Section 3.2 does not use the coupling with imaginary geometry. We derive it directly using a refined analysis of the Radon–Nikodym derivative $M_{\mathbf{t}}(\mathcal{G}_{\mu})$ in Definition 3.6.
@@
$\frac{M_{(t_{1},t_{2})}(\mathcal{G}_{\mu})}{M_{(t_{1},0)}(\mathcal{G}_{\mu})}=$ $\underbrace{\frac{\phi_{\mathbf{t},2}^{\prime}\left(\xi_{t_{2}}^{(2)}\right)^{h}g_{\mathbf{t},2}^{\prime}(0)^{\tilde{h}}\exp\left(\frac{c}{2}m_{\mathbf{t}}\right)}{\left(\phi_{t_{1}}^{(1)}\right)^{\prime}(\theta_{2})^{h}\left(g_{t_{1}}^{(1)}\right)^{\prime}(0)^{\tilde{h}}}}_{=:P_{\mathbf{t}}}\times g_{\mathbf{t},1}^{\prime}(0)^{\frac{3-\mu^{2}}{2\kappa}}\frac{\mathcal{G}_{\mu}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)}{\mathcal{G}_{\mu}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)}\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{h}.$
@@
$g_{\mathbf{t},1}^{\prime}(0)^{\frac{3-\mu^{2}}{2\kappa}}\frac{\mathcal{G}_{\mu}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)}{\mathcal{G}_{\mu}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)}\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{h}.$ (3.16)
Note that the law of radial $\operatorname{SLE}_{\kappa}(2)$ in $\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to 0 with force point $\eta^{(1)}_{t_{1}}$ is the same as radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ weighted by [41]:
$g_{\mathbf{t},1}^{\prime}(0)^{\frac{3}{2\kappa}}\frac{\mathcal{G}_{0}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)}{\mathcal{G}_{0}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)}\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{h}.$ (3.17)
Combining (3.16) and (3.17), we see that the conditional law of $\eta^{(2)}_{[0,t_{2}]}$ given $\eta^{(1)}_{[0,t_{1}]}$ is the same as radial $\operatorname{SLE}_{\kappa}(2)$ in $\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to 0 with force point $\eta^{(1)}_{t_{1}}$ weighted by
$R_{\mathbf{t}}=$ $g_{\mathbf{t},1}^{\prime}(0)^{\frac{-\mu^{2}}{2\kappa}}\frac{\mathcal{G}_{\mu}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)}{\mathcal{G}_{\mu}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)}\frac{\mathcal{G}_{0}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)}{\mathcal{G}_{0}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)}$
$=$ $g_{\mathbf{t},1}^{\prime}(0)^{\frac{-\mu^{2}}{2\kappa}}\exp\left(\frac{\mu}{\kappa}\left(\theta_{\mathbf{t}}^{(1)}+\theta_{\mathbf{t}}^{(2)}\right)-\frac{\mu}{\kappa}\left(\xi_{t_{1}}^{(1)}+\phi_{t_{1}}^{(1)}(\theta_{2})\right)\right).$ (3.18)
We will prove in Lemma 3.10 that the law of radial $\operatorname{SLE}_{\kappa}(2)$ in $\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to 0 with force point $\eta^{(1)}_{t_{1}}$ converges to chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)};\mathrm{e}^{\mathfrak{i}\theta_{2}},0)$ as $t_{1}\to\infty$; and we will show in Lemma 3.11 that $R_{\mathbf{t}}\to 1$ almost surely as $t_{1}\to\infty$. Combining these two parts, the conditional law of $\eta^{(2)}_{[0,t_{2}]}$ given $\eta^{(1)}$ is the same as chordal $\operatorname{SLE}_{\kappa}$ as desired. ∎
@@
Fix $\kappa\in(0,4]$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. We assume $\eta^{(1)}\in\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}};0)$, namely a simple curve in $\mathbb{D}$ from $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ to 0.
  • For $t_{1}\in(0,\infty)$, we denote by $Q_{t_{1}}$ the law of a radial $\operatorname{SLE}_{\kappa}(2)$ in $\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to 0 with force point $\eta^{(1)}_{t_{1}}$.
  • We denote by $Q_{\infty}$ the conditional law of a chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)};\mathrm{e}^{\mathfrak{i}\theta_{2}},0)$.
Suppose $U$ is any neighborhood of $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ in $\mathbb{D}\setminus\eta^{(1)}$ such that there is a positive distance between $U$ and $\eta^{(1)}$ and let $\tau$ be the first time $\eta^{(2)}$ exits $U$. Then the law of the curve $\eta^{(2)}$ restricted to $0,\tau]$ under $Q_{t_{1}}$ and $Q_{\infty}$ are absolutely continuous. We have
$\lim_{t_{1}\to\infty}\frac{\mathrm{d}Q_{t_{1}}}{\mathrm{d}Q_{\infty}}\left(\eta^{(2)}_{[0,\tau]}\right)=1,\quad Q_{\infty}-a.s.$
@@
The convergence of radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ to chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)};\mathrm{e}^{\mathfrak{i}\theta_{2}},0)$ is well known, see e.g. [21, Lem. 3.2] and [26]. In our setup, we need the convergence of radial $\operatorname{SLE}_{\kappa}(2)$ process.
We fix a sequence of conformal maps $f_{t_{1}}:\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}\to\mathbb{D}$ and a conformal map $f_{\infty}:\mathbb{D}\setminus\eta^{(1)}\to\mathbb{D}$ such that
$f_{t_{1}}(\eta^{(1)}_{t_{1}})=\mathrm{e}^{\mathfrak{i}\theta_{1}},\quad f_{t_{1}}(\mathrm{e}^{\mathfrak{i}\theta_{2}})=\mathrm{e}^{\mathfrak{i}\theta_{2}},\qquad f_{\infty}(0)=\mathrm{e}^{\mathfrak{i}\theta_{1}},\quad f_{\infty}(\mathrm{e}^{\mathfrak{i}\theta_{2}})=\mathrm{e}^{\mathfrak{i}\theta_{2}};$
and that $f_{t_{1}}$ converges to $f_{\infty}$ locally uniformly. Note that $(f_{t_{1}})_{*}(Q_{t_{1}})$ is the same as radial $\operatorname{SLE}_{\kappa}(2)$ in $\mathbb{D}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to $w=f_{t_{1}}(0)$ with force point $\mathrm{e}^{\mathfrak{i}\theta_{1}}$, and $(f_{\infty})_{*}(Q_{\infty})$ is the same as chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{2}},\mathrm{e}^{\mathfrak{i}\theta_{1}})$.
The law of radial $\operatorname{SLE}_{\kappa}(2)$ in $\mathbb{D}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to $w$ with force point $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ is the same as the law of radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ weighted by the local martingale [41]:
$g_{t}^{\prime}(0)^{-\tilde{h}}\phi_{t}^{\prime}(\theta_{1})^{h}\left(\sin((\xi_{t}-\phi_{t}(\theta_{1}))/2)\right)^{\frac{2}{\kappa}}$
$\times|g_{t}^{\prime}(w)|^{\tilde{h}+\frac{3}{2\kappa}}(1-|g_{t}(w)|^{2})^{\frac{(\kappa-8)^{2}}{8\kappa}}|g_{t}(w)-\mathrm{e}^{\mathfrak{i}\xi_{t}}|^{1-\frac{8}{\kappa}}|g_{t}(w)-\mathrm{e}^{\mathfrak{i}\phi_{t}(\theta_{1})}|^{1-\frac{8}{\kappa}}.$ (3.19)
The law of chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{2}},\mathrm{e}^{\mathfrak{i}\theta_{1}})$ is the same as the law of radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ weighted by the local martingale [41]:
$g_{t}^{\prime}(0)^{-\tilde{h}}\phi_{t}^{\prime}(\theta_{1})^{h}\left(\sin((\xi_{t}-\phi_{t}(\theta_{1}))/2)\right)^{-2h}.$ (3.20)
Combining (3.3) and (3.20), the law $(f_{t_{1}})_{*}(Q_{t_{1}})$ is the same as $(f_{\infty})_{*}(Q_{\infty})$ weighted by the local martingale
$M_{t}(w)=$ $\left(\frac{\sin((\xi_{t}-\phi_{t}(\theta_{1}))/2)}{\sin((\theta_{2}-\theta_{1})/2)}\right)^{\frac{8}{\kappa}-1}|g_{t}^{\prime}(w)|^{\tilde{h}+\frac{3}{2\kappa}}\left(\frac{1-|g_{t}(w)|^{2}}{1-|w|^{2}}\right)^{\frac{(\kappa-8)^{2}}{8\kappa}}$
$\times\left|\frac{g_{t}(w)-\mathrm{e}^{\mathfrak{i}\xi_{t}}}{w-\mathrm{e}^{\mathfrak{i}\theta_{2}}}\right|^{1-\frac{8}{\kappa}}\left|\frac{g_{t}(w)-\mathrm{e}^{\mathfrak{i}\phi_{t}(\theta_{1})}}{w-\mathrm{e}^{\mathfrak{i}\theta_{1}}}\right|^{1-\frac{8}{\kappa}}.$ (3.21)
As $t_{1}\to\infty$, we have $w=f_{t_{1}}(0)\to\mathrm{e}^{\mathfrak{i}\theta_{1}}$. Thus,
$|g_{t}^{\prime}(w)|\to|g_{t}^{\prime}(\mathrm{e}^{\mathfrak{i}\theta_{1}})|,\quad\frac{1-|g_{t}(w)|^{2}}{1-|w|^{2}}\to|g_{t}^{\prime}(\mathrm{e}^{\mathfrak{i}\theta_{1}})|,$
$\left|\frac{g_{t}(w)-\mathrm{e}^{\mathfrak{i}\xi_{t}}}{w-\mathrm{e}^{\mathfrak{i}\theta_{2}}}\right|\to\frac{\sin((\xi_{t}-\phi_{t}(\theta_{1}))/2)}{\sin((\theta_{2}-\theta_{1})/2)},\quad\left|\frac{g_{t}(w)-\mathrm{e}^{\mathfrak{i}\phi_{t}(\theta_{1})}}{w-\mathrm{e}^{\mathfrak{i}\theta_{1}}}\right|\to|g_{t}^{\prime}(\mathrm{e}^{\mathfrak{i}\theta_{1}})|.$
Consequently, $M_{t}(w)\to 1$ almost surely. These give the desired conclusion. ∎
@@
We write $I_{t_{1}}$ for the arc in $\partial\mathbb{D}$ that is the image of both sides of $\eta^{(1)}_{[0,t_{1}]}$ under the conformal map $g_{t_{1}}^{(1)}$ extended to the boundary, see Figure 5. It is easy to see the harmonic measure of $\partial\mathbb{D}$ seen from 0 of the domain $\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}$ is decreasing to 0 as $t_{1}\to\infty$. Therefore,